Shift report 20011117 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Nov 17 2001 - 16:30:34 EST

  • Next message: Apache: "Shift report 20011118 16:00-24:00"

    ------------------------------------------------------------
    Date: Sat Nov 17 16:30:34 2001, Shift: 08:00-16:00
    Supervisor: James Norris, with: 
    Log book#11 pages 4 - 5
    
    Shift Summary:
    ------------------------------------------------------------
    1. At 08:00 Quench Link Interlock (initial ETA for beam was 09:00)
       At 09:55 beam injection began
       At 10:35 beam stored & cogged
       At 11:25 ZDC=235.7<everyone, called MCR requesting steering.
       At 11:40 noticed that ZDC=331=~STAR ZDC=383
       At 13:10 ZDC=150=~2/3 of initial, and <1/2 of ZDC after steering, 
          and voted with PHOBOS to dump.
       At 14:30 beam dumped.
       At 14:45 beam injection began
       At 16:00 beam stored & cogged
    
    2. Pedestal Run 5855
       TPC Pedestal Run 5857
       DVM Run 5858
       Data Run 5859 - 11:07-11:11 - Trigger 3 not enabled
       Data Run 5860 - 11:12-14:00 - Trigger 3 enabled, NIM Failure, possibly bad C1, H1, H2, and TOFW Data
       Data Run 5864 - 14:01-14:16
    
    3. Anodes were not turned off at beginning of shift even though we did not have beam. For run5860, H1 and H2 did not have TDC start signals, because of the NIM bin power failure (-24V was -15V) and Fan IN/OUT module for TDC start did not work. This was fixed by pulling out a non-used discriminator from the bin. Run5861 seems OK. 
       
    
    3. Anodes were not turned off at beginning of shift even though we did not have beam.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sat Nov 17 2001 - 16:30:53 EST